# Exploring Holomorphic Retracts

### v6

@@ Conclusion
Write $r_{2}=(r_{2}^{(1)},\ldots,r_{2}^{(M)})$. Let $\epsilon:=\min\{r_{2}^{(1)},\ldots,r_{2}^{(M)}\}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $\|v_{2}\|_{\infty}<\epsilon$. Hence $\phi$ maps $D_{1}$ into $D_{2}$. Note that $\phi(p_{1})=p_{2}$ and $\phi^{\prime}(p_{1})(w_{1})=v_{2}g^{\prime}(p_{1})(w_{1})=v_{2}$. Consider the holomorphic map $F:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $F(z,w)=(z,\phi(z))$. Again note that $F(p_{1},w)=(p_{1},p_{2})$ and $F^{\prime}(p)(w_{1},w_{2})=(w_{1},\phi^{\prime}(p_{1})(w_{1}))=(w_{1},v_{2})$. Hence $Z:=\{(z,\phi(z)):z\in D_{1}\}$ is a retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(w_{1},v_{2})\in T_{p}Z$, where $w_{1}\in span_{\mathbb{C}}<v_{1}>$. ∎
Next we shall prove proposition 1.6.
Proof of proposition 1.6. Let us begin by considering the set of boundary points of the form $q/h(q)$ for $q$ varying within $D\setminus A$. Denote the subset of such boundary points satisfying the additional condition that $\partial D$ is non-convex at $q/h(q)$, by $S^{h}$. It follows from the hypotheses that $S^{h}$ is a dense subset of $\partial D$ (as is not difficult to see). Next, we observe that it follows from the above lemma 8.1 together with corollary 2.33, there is no one-dimensional retract through 00 in the direction of $p\in S^{h}$. Note that lemma 8.1 is applicable here owing to the fact that bounded balanced pseudoconvex domains are guaranteed to have a Stein neighbourhood basis by theorem 1.17.16 in [Jrnck_Frst_stps].
With these observations, we are ready to prove the proposition by contradiction. Indeed, if $F$ were a biholomorphism mapping $D\setminus A$ onto a product domain $D$, then through $F(0)$ we have a retract along those directions given by proposition 8.4 – note that the hypothesis about Caratheodory hyperbolicity therein is verified owing to the fact that $D\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D\setminus A$ must be Caratheodory hyperbolic as well. In particular, there always exist a non-trivial retract (necessarily of dimension one here) of $D$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$. By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exist a (non-trivial) retract of $D$ through the origin with its tangent direction at 00 belonging to $S^{\prime}$. But this contradicts the hypotheses that $D$ is not convex at an open dense subset of $\partial D$ which implies that for each non-convex boundary point $a$ of $D$, there are no retracts of $D\setminus A$ passing through the origin in the direction of $a$. This finishes the proof. ∎
@@
Though we had assumed in the statement of proposition 1.6 that the dimension $n=2$, the above proof goes through for arbitrary dimension to show atleast that if $D$ is a balanced domain of holomorphy in $\mathbb{C}^{N}$ with the same assumptions as in that proposition (and $A$ an analytic subset not passing through origin), then $D\setminus A$ is not biholomorphic to a product domain with a one-dimensional factor; in-particular, it is not completely reducible meaning that $D\setminus A$ is not biholomorphic to a product of one-dimensional domains. How about not complete reducibility of the domain but an intermediate one? Establishing such a complete non-existence of a product structure for a large class of domains is rather challenging. Let us only remark that as a direct application of proposition 8.4, we may derive the following corollary from theorem 1.9 proved in an earlier section.
Corollary 8.6.
@@
Let $D_{q}$ denote the $\ell_{q}$-‘ball’ for $10<q<$ 1 and $A$ any analytic subset of $D_{q}$ not containing the origin. Then, neither $D_{q}$ nor $D_{q}\setminus A$, is biholomorphic to a product domain.
@@
We conclude this section with a complete determination of retracts of $D_{h}\setminus A$ through the origin, where $D_{h}$ is the polynomial polyhedron discussed around the beginning of this section. As above, $A$ is an analytic subset of $D_{h}$ not passing through origin, which is allowed to be empty but not $D_{h}$ itself. This of-course is tantamount to determining retracts of $D_{h}$ by lemma 8.1 and we may therefore forget about $A$.
@@
The point here is to record at-least one case of a (non-convex but pseudoconvex) domain with $\mathbb{C}$-extremal but not holomorphically extreme boundary; thereby an example where proposition 1.4 does not apply but the modified form of Vesentini’s result as in theorem 1.3 does. Indeed, $D_{h}$ serves this purpose as we now discuss. First, note that almost every point of $\partial D_{h}$ fails to be holomorphically extreme; specifically, every point of $\partial D_{h}$ which is not in the intersection of the real analytic varieties $H_{1}$ given by $|z^{2}-w^{2}|=1$ and $H_{2}$ given by $|zw|=2$. However, every point of $\partial D_{h}$ is a $\mathbb{C}$-extremal boundary point; this is substantiated by realizing $D_{h}$ as the intersection of the pair of balanced domains $D_{1},D_{2}$ given respectively as the sub-level sets $|z^{2}-w^{2}|<1$ and $|zw|<2$ both of which have (weakly) $\mathbb{C}$-extremal boundaries and then using proposition 3.1.
Now, we contend that for certain special points $p$ (which we shall precisely determine in due course) of $\partial D_{h}$ in the intersection $H_{1}\cap H_{2}$, there exists precisely one retract through the origin in the direction of $p$ which is given by the linear subspace spanned by $p$; along directions given by any other boundary point there is no retract, completing the characterization of retracts through the center of $D_{h}$. We shall presently put forth the proof of this claim. Let us begin by settling at the outset, the non-existence of retracts along directions given by $D_{h}\setminus(H_{1}\cap H_{2})$: this follows by noting that every point of this open piece of the boundary is a non-convex boundary point of $D_{h}$ and invoking corollary 2.33. With this we may restrict ourselves to points of $H_{1}\cap H_{2}\subset\partial D_{h}$ and proceed to discussing retracts through such points alone; it also being borne in mind that by retracts we mean the linear ones for, by theorem 1.3 there are no non-linear holomorphic retracts through the origin in $D_{h}$. To proceed forth indeed then, pick any point $p=(p_{1},p_{2})$ in the intersection $H_{1}\cap H_{2}\subset\partial D_{h}$; so, $|p_{2}^{2}-p_{1}^{2}|=1$ and $|p_{1}p_{2}|=2$. Consider the linear projection $L$ from $\mathbb{C}^{2}$ onto the one-dimensional complex subspace spanned by $p$ denoted $\langle p\rangle$, defined by $L(z,w)=\lambda(z,w)p$, where
@@
$I_{0}=\text{ connected component containing the origin in $\mathbb{R}$ of its subset }\Big{\{}\theta\in\mathbb{R}\;:\;\sin^{2}\theta\leq\frac{1}{16}\Big{\}}.$
@@
to not lose sight of our main goal in the midst of computations, we keep this aside for a moment, and observe that our main goal translates here to checking only whether: $\theta_{0}$ is a point of global maximum for $h$ or not. This is due to the following observation: if at all $p$ is a a point of maximum for $f(z,w)=\tilde{f}(r,\theta)$, then $\theta_{0}$ must be a maximum point for $h$. Now, owing to its smoothness on the open interval $I^{0}$ (the interior of $I$), a point of global maximum for $h$ on $I^{0}$ if any, is not only a point of local maximum but also a point about which $h$ is smooth and so, the standard derivative tests of calculus are applicable. So, if at all $\theta_{0}$ is a point of maximum for $h$ on the compact interval $I$, we must have $h^{\prime}(\theta_{0})=0$ or that $\theta_{0}$ is one of the two boundary points of $I$. But then we infer from (8.14) that: $h^{\prime}(\theta_{0})=0$ implies that $\theta_{0}=0,\pm\pi/2\text{ or }\pi$. Thus as soon as $\theta_{0}$ is not one among the very few possibilities: 00, ±$\pm\pi/2$, $\pi$, or the remaining end-points of the intervals occurring in the definition of $I$, it cannot even be a point of local extremum for $h$. Moreover, the pair of possibilities $\theta_{0}=\pm\pi/2$ are ruled out simply because $h(\theta_{0})$ is not real and therefore not of relevance for our consideration of the function $h$.
We presently proceed to dispose off the cases $\theta_{0}=\pm\epsilon_{0}$ where $\epsilon_{0}$ be the smallest positive number such that $\sin^{2}\epsilon_{0}=1/16$ which means that ±$\pm\epsilon_{0}$ are the end-points of the interval $I$; by showing that $h_{\theta_{0}}(\theta)$ does not attain its (global) maximum on $I$ at $\theta=\theta_{0}=\epsilon_{0}$ – the argument for the other case $-\theta_{0}$ being similar is skipped. Now when $\theta_{0}=\epsilon_{0}$, the expression of $h(\theta)$ simplifies a bit and reads:
@@
Then: firstly, every holomorphic retract through the origin is linear. More specifically and precisely, there are no retracts in the directions of $p\in\partial D_{h}\setminus H$, where $H$ is the real analytic manifold cut out by $|z^{2}-w^{2}|=1$ and $|zw|=2$. As for points $p\in H$, there is a holomorphic retract in the direction of $p$ if and only if it is of the form $(e^{it_{1}}r_{1},e^{it_{2}}r_{2})$ where $r_{1},r_{2}$ are positive real numbers satisfying $r_{1}^{2}-r_{2}^{2}=\pm 1$ and $r_{1}r_{2}=2$ while $t_{1},t_{2}\in\mathbb{R}$ with $|t_{2}-t_{1}|$ being either 00 or $\pi$ (i.e., the arguments of $p_{1},p_{2}$ are either same or differ by $\pi$); and every such retract is linear.
Let us remark that there are of-course points $p$ in $H$ other than those of the form mentioned in the theorem and note more importantly, that those points are precisely the ones in the boundary of the domain $D_{h}$ which are farthest from the origin. Let us briefly justify these claims. For the former, note that for any $\alpha_{0}$ satisfying $160\leq\alpha_{0}\leq 1/$ 16, we can find $p=(p_{1},p_{2})=\left(re^{i\theta_{1}},(2/r)e^{i\theta_{2}}\right)\in H_{1}\cap H_{2}$ such that $\sin^{2}\theta_{0}=\alpha_{0}$, where $\theta_{0}=\theta_{1}-\theta_{2}$. Now choose $\theta_{0}\in I$ such that $\sin^{2}\theta_{0}=\alpha_{0}$. Write $\theta_{0}=\theta_{1}-\theta_{2}$ for some $\theta_{1},\theta_{2}\in[0,2\pi)$. Using this $\theta_{0}$, define

### v7

@@ Conclusion
Write $r_{2}=(r_{2}^{(1)},\ldots,r_{2}^{(M)})$. Let $\epsilon:=\min\{r_{2}^{(1)},\ldots,r_{2}^{(M)}\}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $\|v_{2}\|_{\infty}<\epsilon$. Hence $\phi$ maps $D_{1}$ into $D_{2}$. Note that $\phi(p_{1})=p_{2}$ and $\phi^{\prime}(p_{1})(w_{1})=v_{2}g^{\prime}(p_{1})(w_{1})=v_{2}$. Consider the holomorphic map $F:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $F(z,w)=(z,\phi(z))$. Again note that $F(p_{1},w)=(p_{1},p_{2})$ and $F^{\prime}(p)(w_{1},w_{2})=(w_{1},\phi^{\prime}(p_{1})(w_{1}))=(w_{1},v_{2})$. Hence $Z:=\{(z,\phi(z)):z\in D_{1}\}$ is a retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(w_{1},v_{2})\in T_{p}Z$, where $w_{1}\in\operatorname{span}_{\mathbb{C}}\{v_{1}\}$. ∎
We shall now employ this proposition in applying retracts to establish various biholomorphic inequivalences amongst members of $3$ classes of domains namely, analytic polyhedra, product domains and co-analytic domains i.e., domains of the form $D\setminus A$ wherein $D$ is a domain of holomorphy, an assumption that we generally make unless explicitly stated otherwise. all such domains have the common feature that their boundaries are not entirely smooth and contain non-trivial analytic varieties. On the other hand (as also mentioned in the introductory section), there are differences in the location of the analytic boundary pieces: for instance, in a co-analytic domain there are interior analytic boundary points whereas analytic polyhedra are devoid of such ‘interior’ boundary points, leading them to be biholomorphically inequivalent. To transform this intuition into a rigorous proof as well as to attain a more result a bit more general, we note that co-analytic domains are not hyperconvex whereas analytic polyhedra are. The biholomorphic invariance of hyperconvexity immediately renders the desired inequivalence but take a moment to record this in the following proposition among other related statements in order to be able to refer and make use of it in the sequel. We first formalize the aforementioned notion of boundary points lying in the ‘interior’ of the domain, which we will be used later.
Definition 8.6.
Let $\Omega$ be a domain in $\mathbb{C}^{N}$ and $p\in\partial\Omega$. We say that $p$ is an interior boundary point of $\Omega$ if there exists a neighbourhood $U_{p}$ of $p$ such that $U_{p}\cap\overline{\Omega}=U_{p}$.
Next, we prove a lemma regarding the product of hyperconvex domains in $\mathbb{C}^{N}$.
Lemma 8.7.
Let $D_{1},D_{2}$ be domains in $\mathbb{C}^{N},\mathbb{C}^{M}$ respectively. Then $D:=D_{1}\times D_{2}$ is hyperconvex if and only if $D_{1}$ and $D_{2}$ are hyperconvex.
Proof.
Assume that $D:=D_{1}\times D_{2}$ is hyperconvex and fix a point $(z_{0},w_{0})\in D$. Note that $D_{1}\times\{w_{0}\}$ and $\{z_{0}\}\times D_{2}$ are retracts of $D$. By proposition 2.49, $D_{1}$ and $D_{2}$ are hyperconvex.
Conversely, assume that $D_{1},D_{2}$ are hyperconvex. Let $\varphi_{1},\varphi_{2}$ be psh exhaustion functions of $D_{1},D_{2}$ respectively. Consider the function $\varphi$ on $D_{1}\times D_{2}$ defined by $\varphi(z,w)=\max\{\varphi_{1}(z),\varphi_{2}(w)\}$. Then $\varphi$ is plurisubharmonic and negative on $D_{1}\times D_{2}$. For each $c<0$, let $S^{c}:=\{(z,w)\in D:\varphi(z,w)<c\}$, $S_{1}^{c}:=\{z\in D_{1}:\varphi_{1}(z)<c\}$ and $S_{2}^{c}:=\{w\in D_{2}:\varphi_{2}(w)<c\}$ . Note that
$(z,w)\in S^{c}\Longleftrightarrow\varphi_{1}(z)<c~~\text{and}~\varphi_{2}(w)<c\Longleftrightarrow(z,w)\in S_{1}^{c}\times S_{2}^{c}.$
This shows that $S^{c}=S_{1}^{c}\times S_{2}^{c}$. Since $S_{1}^{c}$ and $S_{2}^{c}$ are relatively compact subsets of $D_{1},D_{2}$ respectively, $S$ is a relatively compact subset of $D$. Hence, $D_{1}\times D_{2}$ is hyperconvex. ∎
Proposition 8.8.
  1. (i)
    Let $D_{1}$ be a complex manifold, $A_{1}$ be a non-empty analytic subset of $D_{1}$ and let $D_{2}$ be a bounded hyperconvex domain in $\mathbb{C}^{N}$, $A_{2}$ be an (possibly empty) analytic subset of $D_{2}$. If $F:D_{1}\setminus A_{1}\to D_{2}\setminus A_{2}$ is a proper holomorphic map then $F$ can be extended as a holomorphic map $\tilde{F}$ from $D_{1}$ to $D_{2}$ with $\tilde{F}(A_{1})\subset A_{2}$. Consequently, there does not exist any proper holomorphic map from $D_{1}\setminus A_{1}$ to $D_{2}$ (as follows by considering the case $A_{2}=\phi$). In particular, there does not exist a biholomorphic map from $D_{1}\setminus A_{1}$ onto $D_{2}$.
  2. (ii)
    Let $D_{1},D_{2}$ be bounded hyperconvex domains in $\mathbb{C}^{N}$ and $A_{1},A_{2}$ be non-empty analytic subsets of $D_{1},D_{2}$ respectively. If $D_{1}\setminus A_{1}$ is biholomorphic to $D_{2}\setminus A_{2}$ then $D_{1}$ is biholomorphic to $D_{2}$.
Proof.
Assume that there exists a proper holomorphic map $F$ mapping $D_{1}\setminus A_{1}$ to $D_{2}\setminus A_{2}$. By applying the Riemann removable singularity theorem to each of the component functions of $F$, we get a holomorphic extension of $F$ which we denote by $\tilde{F}$. The fact that $F$ maps $D_{1}\setminus A_{1}$ into $D_{2}\setminus A_{2}$ implies that $\tilde{F}$ maps $D_{1}$ into $\overline{D_{2}}$. Using the same arguments in the proof of the lemma 8.1, we obtain that $\tilde{F}$ maps $D_{1}$ into $D_{2}$ and $\tilde{F}(A_{1})\subset A_{2}$. Considering the special case $A_{2}=\phi$, we get that there does not exist any proper map from $D_{1}\setminus A_{1}$ to $D_{2}$. This finishes the proof of (i).
To prove (ii), let $F$ be a biholomorphic map from $D_{1}\setminus A_{1}$ onto $D_{2}\setminus A_{2}$ and $G$ be the inverse of $F$. By the arguments similar to those used in (i), we obtain that $\tilde{F}$ maps $D_{1}$ into $D_{2}$ and $\tilde{G}$ maps $D_{2}$ into $D_{1}$. It is easy to see that $\tilde{F}$ and $\tilde{G}$ are biholomorphic maps. To prove this, pick an arbitrary point $z_{0}\in A_{1}$ and a sequence $(z_{n})$ in $D_{1}\setminus A_{1}$ converges to $z_{0}\in A_{1}$. Since $\tilde{F}\equiv F$ on $D_{1}\setminus A_{1}$ and $\tilde{G}\equiv G$ on $D_{2}\setminus A_{2}$, $z_{n}=\tilde{G}\circ\tilde{F}(z_{n})\to\tilde{G}\circ\tilde{F}(z_{0})$. By the uniqueness of limits, $\tilde{G}\circ\tilde{F}(z_{0})=z_{0}$. This shows that $\tilde{G}\circ\tilde{F}$ is the identity map on $D_{1}$. Similarly, $\tilde{F}\circ\tilde{G}$ is the identity map on $D_{2}$. This proves that $\tilde{F}$ is a biholomorphism on $D_{1}$ and its inverse is $\tilde{G}$. Hence $D_{1}$ is biholomorphic to $D_{2}$. ∎
As a corollary of part (i) of the above proposition 8.8 and by above lemma 8.7, we obtain the following.
Corollary 8.9.
Let $D$ be a pseudoconvex domain in $\mathbb{C}^{N}$ and $A$ be a non-empty analytic subset of $D$. Then $D\setminus A$ is not biholomorphic to a product of bounded hyperconvex domains.
In particular, it follows immediately from this corollary that an analytic polyhedron is never biholomorphically equivalent to a co-analytic domain.
@@
In general, it can very well happen that both $D$ and $D\setminus A$ are product (pseudoconvex) domains. Indeed the simplest example showing this is $D=\Delta^{2}$ and $D\setminus A=\Delta\times\Delta^{*}$ obtained by taking $A:=\{(z,w)\in\Delta^{2}:w=0\}$; while there are plenty of other examples, the next proposition shows that within the class of bounded balanced domains of holomorphy in dimension two, this is the only one. This proposition also serves as one of the simplest applications of retracts to establishing biholomorphic inequivalences.
Proposition 8.11.
@@
Let $D$ be any bounded balanced domain of holomorphy in $\mathbb{C}^{2}$ (whose Minkowski functional we denote by $h=h_{D}$), which is not biholomorphic to $\Delta^{2}$. Let $A$ be any proper analytic subset of $D$. Then neither $D$ nor $D\setminus A$ is biholomorphic to a product domain.
Proof.
To prove this proposition by contradiction, assume that $D\setminus A$ is biholomorphic to a product domain $\Omega_{1}\times\Omega_{2}$, where $\Omega_{1},\Omega_{2}$ are domains in $\mathbb{C}$. Let $F$ be a biholomorphism from $D\setminus A$ onto $\Omega_{1}\times\Omega_{2}$ and $G$ be the inverse of $F$. Let $(p_{1},p_{2})\in\Omega_{1}\times\Omega_{2}$. Note that $\Omega_{1}\times\{p_{2}\}$ and $\{p_{1}\}\times\Omega_{2}$ are retracts of $\Omega_{1}\times\Omega_{2}$. Let $G$ be a biholomorphism from $\Omega_{1}\times\Omega_{2}$ onto $D\setminus A$. By proposition 2.4, $\tilde{R_{1}}:=G(\Omega_{1}\times\{p_{2}\})$ and $\tilde{R_{2}}:=G(\{p_{1}\}\times\Omega_{2})$ are retracts of $D\setminus A$. This implies that $\tilde{R_{1}},\tilde{R_{2}}$ are biholomorphic to $\Omega_{1},\Omega_{2}$ respectively. Since $D\setminus A$ is biholomorphic to $\Omega_{1}\times\Omega_{2}$, we obtain that $D\setminus A$ is biholomorphic to $\tilde{R_{1}}\times\tilde{R_{2}}$. By lemma 8.1, there exist one-dimensional retracts $R_{1},R_{2}$ of $D$ such that $\tilde{R_{1}}=R_{1}\cap(D\setminus A)$ and $\tilde{R_{2}}=R_{2}\cap(D\setminus A)$.
If $R_{i}\subset D\setminus A$ for $i=1,2$ then $\tilde{R_{i}}=R_{i}$. Since $R_{i}$ is a simply connected Riemann surface, its biholomorphic image $\Omega_{i}$ is a simply connected domain in $\mathbb{C}$. By Riemann mapping theorem, $R_{i}$ is biholomorphic to $\Delta$. Hence we obtain that $D\setminus A$ is biholomorphic to $\Delta^{2}$, which contradicts part (i) of the above proposition 8.8. Next, we consider the case that $R_{i}\cap A\neq\phi$ for at least one $i$. Let $S$ be the set of all points of intersection of $R_{i}$ with $A$. As $R_{i}\cap A$ does not have any interior point and they are of dimension one, $S$ is a discrete set and thereby can be enumerated as $(p_{j})_{j=1}^{\infty}$. Fix one of the $p_{j}$’s say $p_{1}$ for simplicity. There exists an open set $U_{1}\subset R_{i}$, which is contained within at least one chart about $p_{1}$ for $R_{i}$. Let $\phi$ be the biholomorphism from $\tilde{R_{i}}$ onto $\Omega_{i}$. By Casorati-Weiestrass theorem, $p_{1}$ is not an essential singularity for $\phi_{|_{U_{1}^{*}}}$, where $U_{1}^{*}=U_{1}\setminus\{p_{1}\}$. So, the isolated singularity for $\phi$ at $p_{1}$ is removable when $\phi$ is viewed as a map into $\mathbb{C}_{\infty}$. Let $V_{1}:=\phi(U_{1})$ and $V_{1}^{*}=V_{1}\setminus\{q_{1}\}$, where $q_{1}=\phi(p_{1})$. By local-normal form theorem, after passing to suitable charts, we obtain that $\partial(\phi(U_{1}^{*}))=\phi(\partial U_{1})\cup\{q_{1}\}$ with $q_{1}\notin\phi(\partial U_{1})$. Hence $q_{1}$ is an interior boundary point of $V_{1}$. This argument holds for all $p_{j}$. Since $\phi$ is a biholomorphism on $\tilde{R_{i}}$, $\phi_{|_{S}}$ is one-one. Hence $R_{i}$ is biholomorphic to $\Omega_{i}^{\prime}:=\Omega_{i}\cup\phi(S)$. Since $R_{i}$ is simply connected, $\Omega_{i}^{\prime}$ is simply connected and thereby $R_{i}$ is biholomorphic to $\Delta$ by Riemann mapping theorem. Hence there exist discrete subsets $L_{1},L_{2}\subset\Delta$ (at least one of $L_{i}$ is non-empty) such that $\tilde{R_{1}}$ and $\tilde{R_{2}}$ are biholomorphic to $\Delta\setminus L_{1}$ and $\Delta\setminus L_{2}$ respectively. This implies that $D\setminus A$ is biholomorphic to $(\Delta\setminus L_{1})\times(\Delta\setminus L_{2})$. Write $(\Delta\setminus L_{1})\times(\Delta\setminus L_{2})=\Delta^{2}\setminus S^{\prime}$, where $S^{\prime}=(\Delta\times L_{2})\cup(L_{1}\times\Delta)$. Note that $S^{\prime}$ is a non-empty analytic subset of $\Delta^{2}$. By applying the Riemann removable singularity theorem to each of the component functions of $F$, we get a holomorphic extension of $F$ which we denote by $\tilde{F}$. The fact that $F$ maps $D\setminus A$ to $\Delta^{2}\setminus S^{\prime}$ implies that $\tilde{F}$ maps $D$ into $\overline{\Delta}^{2}$. By theorem 1.17.16 in [Jrnck_Frst_stps], both $D$ and $\Delta^{2}$ admit Stein neighborhood bases. Applying the same argument in the proof of the lemma 8.1 and its remark, we conclude that $\tilde{F}$ maps $D$ into $\Delta^{2}$ and $\tilde{G}$ maps $\Delta^{2}$ into $D$. Using the methods similar to the proof of part (ii) of the above proposition 8.8, we obtain that $\tilde{F}$ and $\tilde{G}$ are biholomorphic maps. This proves that $D$ is biholomorphic to the polydisc $\Delta^{2}$. ∎
Next,we prove the corollary of above proposition 8.5 and theorem 1.8.
Proof of Corollary 1.11: If $F$ is a biholomorphic mapping from $D_{q}\setminus A$ onto a product domain $G=G_{1}\times G_{2}$, then through $F(0)$ we have a retract along those directions given by proposition 8.5 – note that the hypothesis about Caratheodory hyperbolicity therein (for $G$ here) is verified owing to the fact that $D_{q}\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D_{q}\setminus A$ must be Caratheodory hyperbolic as well (refer to theorem 18.2.1 of [Jrncki_invrnt_dst]). In particular, there always exists a non-trivial retract (necessarily of dimension one here) of $D_{q}$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$. By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exist a (non-trivial) retract of $D_{q}$ through the origin with its tangent direction at 0 belonging to $S^{\prime}$. But this contradicts theorem 1.8, which implies that $D_{q}$ does not admit any retract along an open piece of directions at the origin. Hence, $D_{q}\setminus A$ cannot be biholomorphic to a product domain. By the same reasoning, it follows that $D_{q}$ itself is not biholomorphic to a product domain. ∎
@@
We conclude this section with a complete determination of retracts of $D_{h}\setminus A$ through the origin, where $D_{h}$ is the polynomial polyhedron (8.1). As above, $A$ is an analytic subset of $D_{h}$ not passing through the origin, which is allowed to be empty but not $D_{h}$ itself. This of-course is tantamount to determining retracts of $D_{h}$ by lemma 8.1 and we may therefore forget about $A$.
@@
We are yet to determine the linear retracts of $D_{h}$ (the motivation for this has already been mentioned in the prequel). First, note that almost every point of $\partial D_{h}$ fails to be holomorphically extreme; specifically, every point of $\partial D_{h}$ which is not in the intersection of the real analytic varieties $H_{1}$ given by $|z^{2}-w^{2}|=1$ and $H_{2}$ given by $|zw|=2$. However, every point of $\partial D_{h}$ is a $\mathbb{C}$-extremal boundary point; this is substantiated by realizing $D_{h}$ as the intersection of the pair of balanced domains $D_{1},D_{2}$ given respectively as the sub-level sets $|z^{2}-w^{2}|<1$ and $|zw|<2$ both of which have (weakly) $\mathbb{C}$-extremal boundaries and then using proposition 3.1.
Now, we contend that for certain special points $p$ (which we shall precisely determine in due course) of $\partial D_{h}$ in the intersection $H_{1}\cap H_{2}$, there exists precisely one retract through the origin in the direction of $p$ which is given by the linear subspace spanned by $p$; along directions given by any other boundary point there is no retract, completing the characterization of retracts through the center of $D_{h}$. We shall presently put forth the proof of this claim. Let us begin by settling at the outset, the non-existence of retracts along directions given by $D_{h}\setminus(H_{1}\cap H_{2})$: this follows by noting that every point of this open piece of the boundary is a non-convex boundary point of $D_{h}$ and invoking corollary 2.38. With this we may restrict ourselves to points of $H_{1}\cap H_{2}\subset\partial D_{h}$ and proceed to discussing retracts through such points alone; it also being borne in mind that by retracts we mean the linear ones for, by proposition 1.3 there are no non-linear holomorphic retracts through the origin in $D_{h}$. To proceed forth indeed then, pick any point $p=(p_{1},p_{2})$ in the intersection $H_{1}\cap H_{2}\subset\partial D_{h}$; so, $|p_{2}^{2}-p_{1}^{2}|=1$ and $|p_{1}p_{2}|=2$. Consider the linear projection $L$ from $\mathbb{C}^{2}$ onto the one-dimensional complex subspace spanned by $p$ denoted $\langle p\rangle$, defined by $L(z,w)=\lambda(z,w)p$, where
@@
[…]
@@
to not lose sight of our main goal in the midst of computations, we keep this aside for a moment, and observe that our main goal translates here to checking only whether: $\theta_{0}$ is a point of global maximum for $h$ or not. This is due to the following observation: if at all $p$ is a point of maximum for $f(z,w)=\tilde{f}(r,\theta)$, then $\theta_{0}$ must be a maximum point for $h$. Now, owing to its smoothness on the open interval $I^{0}$ (the interior of $I$), a point of global maximum for $h$ on $I^{0}$ if any, is not only a point of local maximum but also a point about which $h$ is smooth and so, the standard derivative tests of calculus are applicable. So, if at all $\theta_{0}$ is a point of maximum for $h$ on the compact interval $I$, we must have $h^{\prime}(\theta_{0})=0$ or that $\theta_{0}$ is one of the two boundary points of $I$. But then we infer from (8.15) that: $h^{\prime}(\theta_{0})=0$ implies that $\theta_{0}=0,\pm\pi/2\text{ or }\pi$. Thus as soon as $\theta_{0}$ is not one among the very few possibilities: 0, $\pm\pi/2$, $\pi$, or the remaining end-points of the intervals occurring in the definition of $I$, it cannot even be a point of local extremum for $h$. Moreover, the pair of possibilities $\theta_{0}=\pm\pi/2$ are ruled out simply because $h(\theta_{0})$ is not real and therefore not of relevance for our consideration of the function $h$.
We presently proceed to dispose off the cases $\theta_{0}=\pm\epsilon_{0}$ where $\epsilon_{0}$ be the smallest positive number such that $\sin^{2}\epsilon_{0}=1/16$ which means that $\pm\epsilon_{0}$ are the end-points of the interval $I$; by showing that $h_{\theta_{0}}(\theta)$ does not attain its (global) maximum on $I$ at $\theta=\theta_{0}=\epsilon_{0}$ – the argument for the other case $\theta_{0}$ being similar is skipped. Now when $\theta_{0}=\epsilon_{0}$, the expression of $h(\theta)$ simplifies a bit and reads:
@@
Then: firstly, every holomorphic retract through the origin is linear. More specifically and precisely, there are no retracts in the directions of $p\in\partial D_{h}\setminus H$, where $H$ is the real analytic manifold cut out by $|z^{2}-w^{2}|=1$ and $|zw|=2$. As for points $p\in H$, there is a holomorphic retract in the direction of $p$ if and only if it is of the form $(e^{it_{1}}r_{1},e^{it_{2}}r_{2})$ where $r_{1},r_{2}$ are positive real numbers satisfying $r_{1}^{2}-r_{2}^{2}=\pm 1$ and $r_{1}r_{2}=2$ while $t_{1},t_{2}\in\mathbb{R}$ with $|t_{2}-t_{1}|$ being either 0 or $\pi$ (i.e., the arguments of $p_{1},p_{2}$ are either same or differ by $\pi$); and every such retract is linear.
Let us remark that there are of-course points $p$ in $H$ other than those of the form mentioned in the theorem and note more importantly, that those points are precisely the ones in the boundary of the domain $D_{h}$ which are farthest from the origin. Let us briefly justify these claims. For the former, note that for any $\alpha_{0}$ satisfying $0\leq\alpha_{0}\leq 1/16$, we can find $p=(p_{1},p_{2})=\left(re^{i\theta_{1}},(2/r)e^{i\theta_{2}}\right)\in H_{1}\cap H_{2}$ such that $\sin^{2}\theta_{0}=\alpha_{0}$, where $\theta_{0}=\theta_{1}-\theta_{2}$. Now choose $\theta_{0}\in I$ such that $\sin^{2}\theta_{0}=\alpha_{0}$. Write $\theta_{0}=\theta_{1}-\theta_{2}$ for some $\theta_{1},\theta_{2}\in[0,2\pi)$. Using this $\theta_{0}$, define